OpenStudy (anonymous):

3-3/s5^-2 simlify please

OpenStudy (tkhunny):

Unable to translate. Can you ad parentheses to clarify intent?

OpenStudy (anonymous):

is it 3 - (3/s5^2) or (3-3)/s5^-2?

OpenStudy (anonymous):

well its 3r/s^-2 r=-3 s=5

OpenStudy (tkhunny):

\(\dfrac{3r}{s^{-2}} = 3rs^2\) -- That's quite a bit easier. Now substitute the given values.

OpenStudy (anonymous):

You just need to substitute the numbers for the letters 3(-3)/5^-2 = -9 x (25) = ____

OpenStudy (anonymous):

oh i understand now! thank you both! idk who to medal though D: you were both awesome

OpenStudy (tkhunny):

Roll the dice and pick Ivan. :-)

OpenStudy (anonymous):

ok haha :)